# Multi-Layer Cycle Benchmarking for high-accuracy error characterization

### v1

@@ Abstract
Accurate noise characterization is essential for reliable quantum computation. Effective Pauli noise models have emerged as powerful tools, offering detailed description of the error processes with a manageable number of parameters, which guarantees the scalability of the characterization procedure. However, a fundamental limitation in the learnability of Pauli fidelities impedes full high-accuracy characterization of both general and effective Pauli noise, thereby restricting e.g., the performance of noise-aware error mitigation techniques. We introduce Multi-Layer Cycle Benchmarking (MLCB), an enhanced characterization protocol that improves the learnability associated with effective Pauli noise models by jointly analyzing multiple layers of Clifford gates. We show a simple experimental implementation and demonstrate that, in realistic scenarios, MLCB can reduce unlearnable noise degrees of freedom by up to $75\%$, improving the accuracy of sparse Pauli-Lindblad noise models and boosting the performance of error mitigation techniques like probabilistic error cancellation. Our results highlight MLCB as a scalable, practical tool for precise noise characterization and improved quantum computation.
@@ I Introduction
Recently, considerable attention has been devoted to characterization schemes that lie between these extremes, offering greater detail than holistic techniques while still remaining scalable [8, 9, 10, 11, 12]. This trend is also driven by the need to implement noise-aware error mitigation techniques [13, 8, 14, 15], which require detailed knowledge of the noise associated with the parallel execution of instructions on large QPUs. Achieving this without incurring the prohibitively high sampling complexity of GST typically requires implementing averaging techniques to recast the original noise into more manageable forms. Equally important is the development of simpler, effective noise models that capture the most relevant features of the noise on the whole QPU while featuring a manageable number of parameters [8, 16, 11].
@@
A prominent example of noise recasting techniques is Randomized Compiling (RC) applied to Clifford gates, which can approximately average, or twirl, noise into so-called Pauli noise [17, 18]. The latter is a type of stochastic noise modeled by Pauli channels, where only errors associated to any Pauli operator occur with a given probability. Incidentally, this makes RC also a valuable error suppression tool, as it gets rid of a large portion of detrimental, so-called coherent errors that get twirled into less detrimental, incoherent ones. This highlights the appeal and practical significance of characterizing Pauli noise [8, 16]. Its relevance is further amplified by the possibility to develop and use effective and scalable versions of Pauli noise, where reasonable assumptions about noise locality can significantly reduce the number of parameters to be determined [8, 9, 16].
@@
Despite these nice features, the characterization of gate-dependent Pauli noise remains fundamentally limited by Pauli learnability constraints, as demonstrated in Ref [19] and confirmed experimentally [8, 20, 17]. As in GST [2, 21], gauge freedom introduces ambiguities, making it impossible to determine certain parameters of Pauli channels independently from State Preparation and Measurement (SPAM) errors. Specifically, Ref. [19] shows that for a set of Clifford gates, the number of these unlearnable noise Degrees of Freedom (DOF) in Pauli channels scales exponentially with the number of qubits $n$. In this context, Cycle Benchmarking (CB) has emerged as an convenient method for accurately determining all learnable DOF of Pauli channels [19].
@@
In this work, we specifically consider the characterization of scalable and effective Pauli noise models and present a enhanced version of CB called Multi-Layer Cycle Benchmarking (MLCB), which significantly increases the number of learnable DOF by leveraging the constraints inherent in the effective noise model. Unlike standard CB implementations [8, 16], MLCB simultaneously analyzes multiple sets of parallel Clifford gates, rather than characterizing each set in isolation. After providing a general framework for implementing MLCB, we focus on the characterization of the IQM GarnetTM [22] quantum processor, which features $20$ superconducting qubits in a square topology. Our results show that MLCB can achieve a remarkable $75\%$ reduction in the number of unlearnable DOF in practically relevant scenarios. Moreover, we experimentally highlight how MLCB-enabled learnable DOF differ significantly from estimates derived through conventional low-accuracy approaches. Additionally, we numerically show how this enhanced learnability offered by MLCB translates into more accurate characterization of effective noise models, focusing on the Sparse Pauli-Lindblad (SPL) noise model introduced in Ref. [8] and widely used to inform noise-aware error mitigation strategies [8, 14, 15]. In this respect, we also investigate numerically how MLCB reduces residual biases in Probabilistic Error Cancellation (PEC), a prominent error mitigation technique that requires precise noise characterization to achieve optimal results [13, 8, 23].
@@
The rest of the paper is organized as follows: Section II provides an overview of the CB technique for Clifford gate layer characterization and discusses Pauli noise learnability constraints. Section III introduces scalable Pauli noise models, with a focus on the SPL model. The MLCB framework is presented in Section IV, followed by its experimental testing on the IQM GarnetTM platform in Section V. Finally, Sections VI and VII illustrate how MLCB enhances characterization accuracy and improves the performance of noise-aware error mitigation techniques.
@@ II Characterization of a Clifford layer
In general, a quantum circuit consists of a list of instructions that can be executed by the quantum hardware, such as quantum gates, reset or state preparation and measurements. The set of qubits targeted by a specific instruction is known as the instruction’s support. We define a layer as a set of instructions featuring non-overlapping supports, which is executed in parallel (although not necessarily simultaneously)  [^1] In the literature, a layer is sometimes also referred to as a cycle [16]. This allows us to conveniently express a generic circuit in terms of a sequence of layers, executed at given time steps. In practically relevant scenarios, the support of native instructions often consists of either one or two qubits. Entangling two-qubit gates are typically Clifford operations, such as iSWAP, CNOT and CZ (the latter being native on IQM’s hardware). When executed on the QPU, two-qubit (2Q) gates are generally noisier than general single-qubit (1Q) rotations. Moreover, because of spatial crosstalk, the execution of a given gate can affect the behavior of neighbouring qubits. Characterizing the noise associated with the execution of a whole Clifford layer $C$, consisting of several 2Q gates, is thus crucial for assessing the errors impacting practical applications. In the following, we introduce the necessary notation by defining the concepts of Pauli twirling, Pauli maps, and the associated Pauli fidelities. We then provide a concise overview of the standard CB protocol [25], which enables the high-accuracy determination of all learnable Pauli fidelities associated with general Pauli noise [19]. This protocol serves as the foundation for the development of our proposed MLCB technique.
II.1 Determining the Pauli fidelities
The ideal implementation of $C$ on a $n$-qubit QPU is represented by a unitary operator $U_{C}\in\mathbb{U}(2^{n})$, corresponding to the ideal quantum channel $\phi(U_{C})$, defined as $\phi(U_{C})[\rho]:=U_{C}\rho U_{C}^{\dagger}$. However, in practice, assuming noise being context-independent and probability-preserving, the actual execution of the layer can be modelled as $\mu(U_{C}):=\phi(U_{C})\circ\tilde{\Lambda}_{C}$, where the completely positive and trace-preserving (CPTP) map $\tilde{\Lambda}_{C}$ captures the effects of the noise. To simplify the form of the noise channel $\tilde{\Lambda}_{C}$ and enable efficient noise characterization, a key technique is twirling. Formally, the twirling of the map is performed by introducing additional layers, before and after the noisy Clifford layer $\phi(U_{C})$, consisting of 1Q operators randomly sampled from the twirling set $\mathbb{T}$. Unless explicitly stated otherwise, in the following we will focus on the standard Pauli twirling, with $\mathbb{T}$ consisting of all possible product combination of 1Q Pauli operators, i.e. $\mathbb{T}=\mathbb{P}_{n}:=\{I,X,Y,Z\}^{\otimes n}$. Instead of executing $U_{C}$, we first implement a twirling layer $T\in\mathbb{T}$, followed by $U_{C}$ and then by the conjugated layer $T_{C}:=U_{C}TU_{C}^{\dagger}\in\mathbb{T}$, which also belongs to $\mathbb{T}$ since the Pauli group is preserved under Clifford conjugation. If we assume that the noise map $\Omega$ associated with the execution of a layer $T$ of 1Q gates is gate-independent, i.e. $\nu(T):=\Omega\circ\phi(T)$ $\forall T$, one can show that uniformly sampling $T$ from the twirling set $\mathbb{T}$ and subsequently averaging the results, denoted as $\langle\cdot\rangle_{T\in\mathbb{T}}$, leads to the following effective evolution
where
is an effective Pauli channel associated with the execution of the Pauli-dressed Clifford layer $C$. The channel $\Lambda_{C}$ corresponds to the noise of the original Clifford layer $C$, which contributes the majority of the noise, together with the noise of the twirling layer $T$. Note that, even if the noise associated with the latter has some gate-dependent components, the whole construction still holds up to quadratic corrections in the deviations from the gate-independent assumption [16].
@@ II.1 Determining the Pauli fidelities
We can now focus on characterizing the Pauli map $\Lambda_{C}$: since it is a linear map, it is straightforward to see that every Pauli operator $P_{\beta}\in\mathbb{P}_{n}$ is an eigenvector of the map, i.e., $\Lambda_{C}[P_{\beta}]=f_{\beta}P_{\beta}$ $\forall P_{\beta}\in\mathbb{P}_{n}$, with the associated eigenvalue $f_{\beta}$. Accurately characterizing the eigenvalues $f_{\beta}$, also known as Pauli fidelities, is one of the main objectives of this paper. Indeed, a perfect knowledge of all the Pauli fidelities $\{f_{i}\}$ allows to determine the elements $\{p_{i}\}$ of the probability distribution in Eq. (2) via the Walsh-Hadamard transform $p_{\alpha}=2^{-n}\sum_{\beta}f_{\beta}(-1)^{\langle\alpha,\beta\rangle_{sp}}$, with $\langle\alpha,\beta\rangle_{sp}$ being the symplectic inner product of Pauli operators $P_{\alpha}$ and $P_{\beta}$, which is zero if they commute and one otherwise.
If it was possible to implement the noise map $\Lambda_{C}$ without the associated unitary $\phi(U_{C})$, the generic Pauli fidelity $f_{\alpha}$ could be measured by implementing a simplified version of the CB protocol, consisting of the following steps: (i) prepare a $+1$ eigenstate of the operator $P_{\alpha}$, (ii) apply $\Lambda_{C}$ a number of times $d$ and, (iii) measure the expectation value of $P_{\alpha}$. Indeed, for $d=1$, one obtains
$\begin{split}&\text{Tr}\left\{P_{\alpha}\Lambda_{C}[\rho_{\alpha}]\right\}=\frac{\text{Tr}\left\{P_{\alpha}\Lambda_{C}[P_{\alpha}]\right\}}{2^{n}}=f_{\alpha}.\end{split}$ (3)
For generic $d$, in the presence of SPAM errors and by using readout twirling [26, 27], one can show that the expectation value is given by $A_{\alpha}(f_{\alpha})^{d}$  [^2] Without readout twirling one would generally get a constant shift $A_{\alpha}(f_{\alpha})^{d}+B_{\alpha}$ By performing this procedure for different values of $d$ and fitting the results to an exponential curve, one can determine $f_{\alpha}$ with high accuracy. Here, high accuracy signifies that the quantity can be measured with multiplicative precision in a SPAM-robust manner, as the SPAM coefficients $A_{\alpha}$ are independent of $d$ [25].
However, in practice, it is not feasible to apply the map $\Lambda_{C}$ in isolation, as the entire entire dressed Clifford layer $C$ needs to be implemented. The resulting characterization protocol becomes what is known as standard CB and the learnability of the Pauli map $\Lambda_{C}$ becomes limited, as we will discuss in the following.
@@
Figure 1: Illustration of the standard CB protocol. In the state preparation stage (left brown area) an eigenstate of a Pauli operator is prepared. The Clifford layer $C$ subject to characterization (in this case, consisting of several parallel CZ gates shown in blue) is then implemented $d$ times, alternating it with layers of 1Q gates (white boxes) which are needed to implement RC. This can be interpreted as the implementation of $d$ dressed Clifford layers. After a final layer of 1Q gates which ensures that RC does not modify the unitary associated with the circuit, the measurement stage (right brown box), which also incorporate readout-twirling, provide us with the expectation value of a Pauli operator (typically the same operator as considered in the state preparation stage).
II.2 Limited learnability of Pauli noise
The standard CB protocol is illustrated in Fig. 1 where, after a state preparation stage (left brown region), the twirled layer $C$ is implemented $d$ times and followed by a final measurement stage (right brown region). Since adjacent 1Q layers can be compiled together [18], the central region of the circuit consists of $d$ instances of $C$ (consisting of parallel 2Q gates shown in blue) which alternate with $d+1$ 1Q gate layers (white boxes). By preparing a $+1$ eigenstate of a Pauli operator $P_{\alpha}$ and measuring its expectation value (using readout twirling), the measurement outcome $\mathsf{O}_{d}^{\alpha}$ reads
$\mathsf{O}_{d}^{\alpha}\propto 2^{-n}\;\text{Tr}\left\{P_{\alpha}\,\Omega\circ\left(\phi(U_{c})\circ\Lambda_{C}\right)^{\circ d}[P_{\alpha}]\right\}.$ (4)
The extra noise map $\Omega$, originating from the last “unpaired” layer of 1Q gates, is not problematic since it can be effectively considered as part of the SPAM error. By contrast, the presence of $\phi(U_{c})$ adds one important complication with respect to the simple scenario discussed before. Indeed, while the Clifford nature of $C$ assures that a Pauli operator is mapped to another Pauli $\phi(U_{C})[P_{\alpha}]=U_{C}P_{\alpha}U_{C}^{\dagger}=P_{\beta}$, there is no guarantee that $P_{\alpha}=P_{\beta}$. This brings us to the definition of the orbit $\mathcal{R}^{C}_{\alpha}$ of a Pauli operator $P_{\alpha}$ as the set of unique Pauli operators that can be reached via the repeated application of the Clifford layer $C$, i.e. $\mathcal{R}^{C}_{\alpha}=\{U_{C}^{j}P_{\alpha}(U_{C}^{\dagger})^{j}\}|j\in\mathbb{N}\}$ [16]. Note that the cardinality of each orbit is upper bounded by $|\mathcal{R}^{C}_{\alpha}|\leq m$, where $m$ is the smallest positive integer such that $(U_{C})^{m}=I$, i.e., the order of Clifford group element $U_{C}$. If $|\mathcal{R}^{C}_{\alpha}|=1$, then the CB experiment described in Eq. (4) results in $\mathsf{O}_{d}^{\alpha}=A_{\alpha}(f_{\alpha})^{d}$. It is thus possible to determine the individual fidelity $f_{\alpha}$ with high accuracy. However, if $|\mathcal{R}^{C}_{\alpha}|>1$, only the product of the corresponding fidelities can be accessed with high accuracy. Indeed, considering $d=k|\mathcal{R}^{C}_{\alpha}|$ (with $k\in\mathbb{N}$), the resulting measurement outcome of the CB experiment is given by the product of the Pauli fidelities, or more explicitly
$\mathsf{O}_{d}^{\alpha}=A_{\alpha}\left(\prod_{\beta\in\{\gamma|P_{\gamma}\in\mathcal{R}^{C}_{\alpha}\}}f_{\beta}\right)^{k}.$ (5)
Importantly, 1Q Clifford gates, such as $S=\sqrt{Z}$ and $\sqrt{X}$, can map each of the three 1Q Pauli operators ($X$, $Y$, and $Z$) into either of the other two. As a result, if a given orbit contains different Pauli operators that share same support, its cardinality can be reduced, thus enabling the determination of more quantities with high accuracy. The combination of this approach, known as interleaved CB, with the standard CB protocol allows to access all the learnable DOF of general Pauli noise [19]. By contrast, the cardinality of orbits consisting of operators with different support cannot be reduced further. Their existence is thus associated with the presence of unlearnable DOF, which can be also understood as a consequence of a fundamental gauge freedom, similar to the one encountered in the context of GST [19, 2]. Specifically, out of the $14^{n}-$ 1 DOF of general Pauli noise, the fundamentally unlearnable ones amount to $2^{n-c}$, where where $c$ is the number of connected components of the pattern transfer graph obtained from the analysis of the orbits (see Ref. [19] for more details).
@@ II.2 Limited learnability of Pauli noise
$P_{\alpha}$ IX IY IZ XI XX XY XZ YI YX YY YZ ZI ZX ZY ZZ
$U_{C}P_{\alpha}U_{C}^{\dagger}$ ZX ZY IZ XZ YY YX XI YZ XY XX YI ZI IX IY ZZ
$\bar{U}_{C}P_{\alpha}\bar{U}_{C}^{\dagger}$ ZY ZX IZ YZ XX XY YI XZ YX YY XI ZI IY IX ZZ
For the sake of concreteness, it is convenient to discuss a simple example: the case of a single CZ gate performed on a $2$-qubit device. In this case, $U_{C}^{2}=I$ limits the cardinality of the orbits to $2$, as it emerges also from Tab. 1, which shows the conjugation of all non-trivial Pauli operators in $\mathbb{P}_{2}$ under a single CZ . Allowing for the usage of additional interleaved $S$ gates, which map $X$ into $Y$ and vice versa, there are $7$ Pauli fidelities which are directly learnable (as they are associated with orbits with cardinality $1$). The remaining $8$ fidelities are not directly measurable but the CB protocols allow to determine, with high accuracy, $8$ products of fidelity pairs. The latter are not all independent from each other since, for example, $f_{IY}f_{ZX}=(f_{IY}f_{ZY})(f_{IX}f_{ZX})/(f_{IX}f_{ZY})$. Still, they allow to impose a total of $6$ high-accuracy constraints on the parameters of the noise model, leaving us with only $2$ unlearnable DOF, which can be identified, for example, with fidelities $f_{IX}$ and $f_{XI}$ (this identification is clearly not unique). More details concerning this simple example are provided in Appendix A.
We conclude this section by emphasizing that the limited learnability of Pauli noise—stemming from the inherent challenge of independently characterizing state preparation and measurement errors [19]—also impacts alternative protocols developed to measure Pauli fidelities. Examples include tensor-network-based noise learning [11] and Averaged Circuit Eigenvalue Sampling (ACES) [9, 12, 10]. For instance, the ACES protocol can reconstruct all relevant fidelities associated with a local Pauli noise model only under the assumption of perfect state preparation [9, 12].

### v2

@@ Abstract
Accurate noise characterization is essential for reliable quantum computation. Effective Pauli noise models have emerged as powerful tools, offering a detailed description of the error processes with a manageable number of parameters, which guarantees the scalability of the characterization procedure. However, a fundamental limitation in the learnability of Pauli eigenvalues impedes a full high-accuracy characterization of both general and effective Pauli noise models, thereby restricting e.g., the performance of noise-aware error mitigation techniques. We introduce Multi-Layer Cycle Benchmarking (MLCB), an enhanced characterization protocol that improves the learnability associated with effective Pauli noise models by jointly analyzing multiple layers of Clifford gates. We show a simple experimental implementation and demonstrate that, in realistic scenarios, MLCB can reduce unlearnable noise degrees of freedom by up to $75\%$, improving the accuracy of sparse Pauli-Lindblad noise models and boosting the performance of error mitigation techniques like probabilistic error cancellation. Our results highlight MLCB as a scalable, practical tool for precise noise characterization and improved quantum computation.
@@ I Introduction
Recently, considerable attention has been devoted to characterization schemes that lie between these extremes, offering greater detail than holistic techniques while still remaining scalable [46, 20, 40, 33, 29]. This trend is also driven by the need to implement noise-aware error mitigation techniques [43, 46, 16, 32] and noise-aware decoding in the context of quantum error correction [30, 44], both requiring detailed knowledge of the noise associated with the parallel execution of instructions on large QPUs. Achieving this without incurring the prohibitively high sampling complexity of GST typically requires implementing averaging techniques to recast the original noise into more manageable forms. Equally important is the development of simpler, effective noise models that capture the most relevant features of the noise on the entire QPU while featuring a manageable number of parameters [46, 7, 33].
@@
A prominent example of noise recasting techniques is Pauli Twirling (PT) applied to Clifford gates, which twirls noise into Pauli noise [31, 48, 26]. The latter is a type of stochastic noise modeled by Pauli channels, where only errors associated with any Pauli operator occur with a given probability. Incidentally, this makes PT also a valuable error suppression tool, as it gets rid of a large portion of detrimental, coherent errors that get twirled into less detrimental, incoherent ones. This highlights the appeal and practical significance of characterizing Pauli noise [46, 7]. Its relevance is further amplified by the possibility to develop and use effective and scalable versions of Pauli noise, where reasonable assumptions about noise locality can significantly reduce the number of parameters to be determined [46, 20, 7].
@@
Despite these advantages, the characterization of gate-dependent Pauli noise remains fundamentally limited by Pauli learnability constraints, as demonstrated in Ref. [9] and confirmed experimentally [46, 15, 26]. As in GST [38, 37], gauge freedom introduces ambiguities, making it impossible to determine certain parameters of Pauli channels independently from State Preparation and Measurement (SPAM) errors. Specifically, Ref. [9] shows that for a set of Clifford gates, the number of these unlearnable noise Degrees of Freedom (DOF) in Pauli channels scales exponentially with the number of qubits $n$. In this context, procedures based on Cycle Benchmarking (CB) [14, 19] have emerged as convenient methods for accurately determining all learnable DOF of Pauli channels [9].
@@
In this work, we specifically consider the characterization of scalable and effective Pauli noise models and present an enhanced version of CB called Multi-Layer Cycle Benchmarking (MLCB), which significantly increases the number of learnable DOF by leveraging the constraints inherent in the effective noise model. Unlike standard CB implementations [46, 7], MLCB simultaneously analyzes multiple sets of parallel Clifford gates, rather than characterizing each set in isolation. After providing a general framework for implementing MLCB, we focus on the characterization of the IQM GarnetTM [1] quantum processor, which features $20$ superconducting qubits in a square topology. Our results show that MLCB can achieve a remarkable $75\%$ reduction in the number of unlearnable DOF in practically relevant scenarios. Moreover, we experimentally highlight how MLCB-enabled learnable DOF differ significantly from estimates derived through conventional low-accuracy approaches. Additionally, we numerically show how this enhanced learnability offered by MLCB translates into more accurate characterization of effective noise models, focusing on the Sparse Pauli-Lindblad (SPL) noise model introduced in Ref. [46] and widely used to inform noise-aware error mitigation strategies [46, 16, 32, 15]. In this respect, we also investigate numerically how MLCB reduces residual biases in Probabilistic Error Cancellation (PEC), a prominent error mitigation technique that requires precise noise characterization to achieve optimal results [43, 46, 22, 15].
@@
The rest of the paper is organized as follows: Sections II and III.1 provide a short overview of the established CB technique for Clifford gate layer characterization, discuss Pauli noise learnability constraints and briefly review scalable Pauli noise models, with a focus on the SPL model. The novel MLCB framework is presented in Section IV, followed by its experimental testing on the IQM GarnetTM platform in Section V. Finally, Sections VI and VII numerically illustrate how MLCB enhances characterization accuracy and improves the performance of noise-aware error mitigation techniques.
@@ II Characterization of a Clifford layer
A quantum circuit can be decomposed into a sequence of layers, where we define a layer as a set of instructions with non-overlapping qubit supports executed in parallel. Characterizing the performance of entire layers, rather than just individual gates, is crucial for assessing errors in practical applications, especially due to effects like spatial crosstalk where operations on some qubits also affect their neighbors.
This work focuses on characterizing the noise of a $n$-qubit Clifford layer $C$ consisting of two-qubit entangling gates, whose execution typically represents the main source of errors. We model the noisy implementation of the layer $C$ as a map $\mu(U_{C})[\rho]:=(\phi(U_{C})\circ\tilde{\Lambda}_{C})[\rho]$, where $\phi(U_{C})[\rho]:=U_{C}\rho U_{C}^{\dagger}$ is associated with the unitary operator $U_{C}\in\mathbb{U}(2^{n})$ and $\tilde{\Lambda}_{C}$ is a completely positive and trace-preserving noisy map. The latter can be simplified into an effective Pauli channel
$\Lambda_{C}[\rho]=\sum_{\alpha}p_{\alpha}P_{\alpha}\rho P_{\alpha}^{\dagger}\quad(P_{\alpha}\in\mathbb{P}_{n}=\{I,X,Y,Z\}^{\otimes n}),$ (1)
by using PT, which requires sandwiching the layer of interest $C$ with two layers of single-qubit Pauli gates [31, 48, 7, 26]. As adjacent layers of single-qubit Pauli operators can then be compiled together [48], the Pauli map $\Lambda_{C}[\rho]$ can be thus understood as the noise affecting the dressed Clifford layer $C$, i.e. the layer $C$ combined with a single layer of (random) Pauli gates. For more details about the twirling procedure, we refer the reader to App. A as well as to Refs. [31, 48, 7].
@@
The task of characterizing the map $\Lambda_{C}[\rho]$ reduces to determining its Pauli eigenvalues $f_{\beta}$, associated with each Pauli operator $P_{\beta}$ as $\Lambda_{C}[P_{\beta}]=f_{\beta}P_{\beta}$ $\forall P_{\beta}\in\mathbb{P}_{n}$. Indeed, these Pauli eigenvalues, sometimes also referred to as Pauli fidelities [45, 9, 7], can be directly related to the error probabilities in Eq. (1) via the Walsh-Hadamard transform $p_{\alpha}=2^{-n}\sum_{\beta}f_{\beta}(-1)^{\langle\alpha,\beta\rangle_{sp}}$, with $\langle\alpha,\beta\rangle_{sp}$ being the symplectic inner product of Pauli operators $P_{\alpha}$ and $P_{\beta}$, which is zero if they commute and one otherwise.
@@
Figure 1: Standard CB protocol: In the state preparation stage (left brown area) an eigenstate of a Pauli operator is prepared. The Clifford layer $C$ subject to characterization (in this case, consisting of several parallel CZgates shown in blue) is then implemented $d$ times, alternating it with layers of single-qubit gates (white boxes) which are needed to implement RC. This can be interpreted as the implementation of $d$ dressed Clifford layers. After a final layer of single-qubit gates which ensures that RC does not modify the unitary associated with the circuit, the measurement stage (right brown box), which also incorporates readout-twirling, provide us with the expectation value of a Pauli operator (typically the same operator as considered in the state preparation stage).
Products of Pauli eigenvalues can be efficiently measured with high-accuracy [19] using CB-structured circuits [14], as sketched in Fig. 1. At first, a +1+1 eigenstate of the Pauli operator $P_{\alpha}$ is prepared; then the Clifford layer $C$ to be characterized is applied $d=km$ times (with $k,m\in\mathbb{N}_{>0}$); finally, the expectation value of $P_{\alpha}$ is measured. By implementing PT targeted to both gate and readout errors [42, 45, 46], thus properly ensuring stochasticity of the latter [25, 4], one can directly access the quantity
$A_{\alpha}\left(\prod_{\beta|P_{\beta}\in\mathcal{R}_{\alpha}^{C}}f^{C}_{\beta}\right)^{k}.$ (2)
Here, the prefactor $A_{\alpha}$ fully accounts for SPAM errors (as constant offsets are eliminated by readout twirling) and $\mathcal{R}_{\alpha}^{C}$ is the orbit of the operator $P_{\alpha}$ under the application of $C$ [7]. More specifically, we define it as the tuple
$\mathcal{R}_{\alpha}^{C}=(P_{\alpha},\phi(U_{C})[P_{\alpha}],\;\dots\;,\phi(U_{C})^{\circ(m-1)}[P_{\alpha}]),$ (3)
where the orbit’s cardinality $m\geq 1$ is the smallest integer such that $\phi(U_{C})^{\circ(m)}[P_{\alpha}]=P_{\alpha}$. Note that $m$ is upper bounded by the smallest positive integer $u$ such that $(U_{C})^{u}=I$. By repeating the experiment for different values of $k$, it is thus possible to determine with multiplicative precision and in a SPAM-robust way [24, 19] the products of the Pauli eigenvalues associated with the operators belonging to each orbit $\mathcal{R}_{\alpha}^{C}$. Importantly, if an orbit features at least a pair of Pauli operators with the same support, its cardinality can be reduced by interleaving the layer $C$ with additional layers of specific single-qubit Clifford gates, which can be compiled together with the twirling layers [9, 46]. In the remainder of this paper, we refer to this whole procedure as the standard CB protocol for measuring Pauli eigenvalues.
The existence of orbits that cannot be reduced to a single element shows that some Pauli eigenvalues cannot be individually learned with high accuracy. This is not a flaw of the standard CB protocol but rather a manifestation of the general limited learnability of Pauli noise, stemming from the existence of a fundamental gauge freedom akin to the one encountered in the context of GST [9, 38]. Specifically, Ref. [9] shows that out of the $4^{n}-1$ DOF of general Pauli noise, the fundamentally unlearnable ones amount to $2^{n-c}$, where $c$ is the number of connected components of the pattern transfer graph based on the orbits analysis. As such, the limited learnability of Pauli noise also impacts alternative protocols developed to measure Pauli eigenvalues. Examples include tensor-network-based noise learning [33] and Averaged Circuit Eigenvalue Sampling (ACES) [20, 29, 40]. For instance, the ACES protocol can reconstruct all relevant Pauli eigenvalues associated with a local Pauli noise model only by arbitrarily fixing the gauge, e.g. by assuming perfect state preparation [20, 29].
A Pauli eigenvalue $f_{\alpha}$ being unlearnable does not imply that it cannot be estimated. For instance, this can be achieved using a modified CB-structured circuit with $d=1$, where an eigenstate of $P_{\alpha}$ is prepared and the expectation value of the conjugated Pauli operator $P_{\beta}=\phi(U_{C})[P_{\alpha}]$ is measured [35, 7, 46]. Those alternative approaches, however, lack the SPAM-robustness and the multiplicative precision that the standard CB protocol provides for learnable quantities. In what follows, we will collectively refer to strategies used to estimate unlearnable Pauli eigenvalues as low-accuracy protocols, contrasted with the high accuracy of protocols like standard CB.
@@
$P_{\alpha}$ IX IY IZ XI $X$ XY XZ YI YX $Y$ YZ ZI ZX ZY $Z$
$U_{C}P_{\alpha}U_{C}^{\dagger}$ ZX ZY IZ XZ $Y$ YX XI YZ XY $X$ YI ZI IX IY $Z$
$\bar{U}_{C}P_{\alpha}\bar{U}_{C}^{\dagger}$ ZY ZX IZ YZ $X$ XY YI XZ YX $Y$ XI ZI IY IX $Z$
For the sake of concreteness, we conclude this section by discussing a simple example: the case of a single CZgate performed on a $2$-qubit device. In this case, $U_{C}^{2}=I$ limits the cardinality of the orbits to $2$, as it emerges also from Tab. 1, which shows the conjugation of all non-trivial Pauli operators in $\mathbb{P}_{2}$ under a single CZ. Allowing for the usage of additional interleaved $S$ gates, which map $X\leftrightarrow Y$, there are $7$ Pauli eigenvalues which are directly learnable, i.e. associated with orbits with cardinality $1$. The remaining $8$ eigenvalues are not directly measurable but the standard CB protocols allow to determine, with high accuracy, $8$ products of eigenvalue pairs. The latter are not all independent from each other since, for example, $f_{IY}f_{ZX}=(f_{IY}f_{ZY})(f_{IX}f_{ZX})/(f_{IX}f_{ZY})$. Still, they allow to impose a total of $6$ high-accuracy constraints on the parameters of the noise model, leaving us with only $2$ unlearnable DOF, which can be identified, for example, with Pauli eigenvalues $f_{IX}$ and $f_{XI}$. Those can only be estimated using low-accuracy protocols. More details concerning this simple example are provided in Appendix B.